OpenStudy (anonymous):

Is a triangle with sides of length 6 ft, 21 ft, 23 ft a right triangle? Explain.

jimthompson5910 (jim_thompson5910):

if you can show that 6^2 + 21^2 = 23^2 is true, then this triangle is a right triangle.

OpenStudy (anonymous):

No 6^2 + 21^2 is not 23^2

OpenStudy (anonymous):

The rule of right angled triangles is that A^2 + B^2 = C^2. A here is 6, B is 21 and C is 23 but that isn't true.

OpenStudy (anonymous):